OpenStudy (anonymous):

help me factorise x^2+17x+46

OpenStudy (akashdeepdeb):

Use the quadratic formula for the quadratic equation f(x) = x2 + 17x + 46 in the form of [ax2+bx+c] where a = 1, b = 17, c = 46 Quadratic formula = \[\frac{-b\pm \sqrt{b^2-4ac}}{2a}\] Got it? :)

OpenStudy (anonymous):

but i want it in the form (x )(x )

OpenStudy (anonymous):

you have to use the formula. the answers will be in decimals if you dont use the formula and leave it in surd form

OpenStudy (anonymous):

but i want it the other way

OpenStudy (anonymous):



OpenStudy (anonymous):

use the formula. the answer will be something like the surd form but it is plus minus so you can easily make it equal to x. form there you can make it (x+ )(x- ).

OpenStudy (anonymous):

what year are u in cause it6's all too confusing for me

OpenStudy (anonymous):

Are you allowed to answer that the formula is "prime" (i.e., can't be answered in integer terms)? Or, can you answer in the surd form?

OpenStudy (anonymous):